OpenStudy (anonymous):

solve by substitution. 2x+y=4 y=x+1 Help ?o;

OpenStudy (anonymous):

substitute y = x + 1 into the first equation \[\implies 2x + x + 1 = 4\] the rest is algebra

OpenStudy (anonymous):

2x+x+1=4= 2*1+1+1

OpenStudy (anonymous):

Thank-you Both!! Soo would'nt it look like this.. 2(x+1)=4 Right ?

OpenStudy (anonymous):

Yep you got it

OpenStudy (anonymous):

Can you help me

OpenStudy (anonymous):

Alright thanks again! & Suree I can try ?(: